# Coherence, Transport, and Chaos in 1D Bose-Hubbard Model: Disorder vs. Stark Potential

### v1

@@ V Conclusions
Our exact numerical study of the 1D BHM reveals the intricate interplay between thermal fluctuations, a SP, and quenched disorder in shaping quantum coherence and phase behavior. By employing a suite of complementary observables, we examined the complex landscape of quantum phases and transitions arising from their combined effects. In the absence of disorder, our results indicate a QPT near $\tau/U\approx 0.17$, where all coherence indicators including condensate fraction, SF fraction, visibility, and the $\ell_{1}$ norm of coherence, exhibit a sharp increase, indicating the transition from a MI to a SF phase. At finite temperatures, this transition is transformed to a smooth crossover. Interestingly, both the $\ell_{1}$-norm and number fluctuations are significantly enhanced in the intermediate regime, driven by thermal activation of delocalized excited states. This demonstrates that thermal effects can enable coherence mechanisms that are absent in the GS.
@@
Our multi-observable analysis reveals that while both the SP and disorder suppress global coherence, they do so through fundamentally distinct mechanisms, namely, energy gradients in the case of the SP and random scattering due to disorder. These distinct origins lead to unique signatures in correlation-based observables. Notably, the $\ell_{1}$-norm of coherence emerges as a particularly sensitive probe, capable of detecting subtle coherence structures that remain hidden to conventional measures. This sensitivity is especially evident in the BG regime, where thermal delocalization enhances coherence among excited states.
These findings reveal the multiplex interplay among tunneling, interactions, disorder, and temperature in phase transition and quantum properties of the strongly correlated bosonic systems. They hold significant implications for quantum simulation platforms based on ultracold atoms. By elucidating how various perturbations reshape the coherence landscape, our study contributes to a deeper understanding of strongly correlated phases beyond the GS. Furthermore, the combined effects of the SP and disorder introduce a richer spectrum of physical phenomena, particularly through the coexistence of deterministic and stochastic localization mechanisms in 1D BH systems.
Acknowledgments: H.A.Z. acknowledges the financial support provided under the postdoctoral fellowship program of P. J. Šafárik University in Košice, Slovakia. H.A.Z. is also supported by the grants of the Slovak Research and Development Agency under contract No. APVV-20-0150 and the Ministry of Education, Research, Development and Youth of the Slovak Republic under contract No. VEGA 1/0695/23. S.H. thanks the School of Particles and Accelerators at the Institute for Research in Fundamental Sciences for their financial support.

### v2

@@ V Conclusions
Our exact numerical study of the 1D BHM reveals the intricate interplay between thermal fluctuations, a SP, and quenched disorder in shaping quantum coherence and phase behavior. By employing a suite of complementary observables, we examined the complex landscape of quantum phases and transitions arising from their combined effects. In the absence of disorder, our results indicate a QPT near $\tau/U\approx 0.17$, where all coherence indicators including condensate fraction, SF fraction, visibility, and the $\ell_{1}$-norm of coherence, exhibit a sharp increase, indicating the transition from a MI to a SF phase. At finite temperatures, this transition is transformed to a smooth crossover. Interestingly, both the $\ell_{1}$-norm and number fluctuations are significantly enhanced in the intermediate regime, driven by thermal activation of delocalized excited states. This demonstrates that thermal effects can enable coherence mechanisms that are absent in the GS.
@@
Our multi-observable analysis reveals that while both the SP and disorder suppress global coherence, they do so through fundamentally distinct mechanisms, namely, energy gradients in the case of the SP and random scattering due to disorder. These distinct origins lead to unique signatures in correlation-based observables. Notably, the $\ell_{1}$-norm of coherence emerges as a particularly sensitive probe, capable of detecting subtle coherence structures hidden from conventional measures. This sensitivity is especially evident in the BG regime, where thermal delocalization enhances coherence among excited states.
These findings reveal the multiplex interplay among tunneling, interactions, disorder, and temperature in phase transition and quantum properties of the strongly correlated bosonic systems. They hold significant implications for quantum simulation platforms based on ultracold atoms. By elucidating how various perturbations reshape the coherence landscape, our study contributes to a deeper understanding of strongly correlated phases beyond the GS. Furthermore, the combined effects of SP and disorder introduce a richer spectrum of physical phenomena, particularly through the coexistence of deterministic and stochastic localization mechanisms in 1D BH systems.
Acknowledgments: H.A.Z. acknowledges the financial support provided under the postdoctoral fellowship program of P. J. Šafárik University in Košice, Slovakia. H.A.Z. is also supported by the grants of the Slovak Research and Development Agency under contract No. APVV-20-0150 and the Ministry of Education, Research, Development and Youth of the Slovak Republic under contract No. VEGA 1/0298/25.